Find the minimum and maximum values of the function 7cosx-24sinx+5

Grade:11

1 Answers

NIHIR SHAH
3 years ago
by using the logik that the maxim and minim value of any trigo fxn a cos x + b sin x is root ( a^2+ b^2) , so 7 cos x + 24 sin x , hence max val = 25 and min = -25 so , min value of function  5-25 = -20 and max val = 5+25 = 30